OpenStudy (anonymous):

How do i do this?! Multiply and simplify: 3x^2-12/5x-10 x 1/2x-4

OpenStudy (anonymous):

\[=3(x ^{2}-4)/5(x-2) * 1/x(x-2) = 3(x-2)(x+2)/5(x-2) *1/2(x-2)\]

OpenStudy (anonymous):

cancel out (x-2) \[3(x+2)/10(x-2)\]